An 1100-kg car traveling at 27.0 m/s starts to slow down and comes to a complete stop in 578 m. What is the magnitude of the average braking force acting on the car?


1
Expert's answer
2021-02-10T18:23:33-0500

"<F>=\frac As=\frac{mv^2}{2s},"

"<F>=\frac{1100\cdot 27^2}{2\cdot 578}=694~N."